# Venus: An Efficient Edge Memory-and-Retrieval System for VLM-based Online Video Understanding

### v1

@@ V-A3 Baselines Methods
    MDF [han2024self] proposes a query-irrelevant method that adopts a self-adaptive strategy to filter the most dominant frames.
@@
    Video-RAG [luo2024video] adopts uniform sampling over video frames and incorporates additional auxiliary data to construct a RAG database that enhances VLM performance.
@@
    AKS [tang2025adaptive] proposes an adaptive keyframe selection method and employs an optimization algorithm to ensure comprehensive coverage of the selected keyframes.
@@
    BOLT [liu2025bolt] designs an inverse transform sampling method to prioritize query-relevant frames, improving performance on multi-source retrieval VQA tasks.
@@ V-C Comparison with Query-Relevant Baselines
We compare Venus with query-relevant baselines across different datasets, VLMs and deployment strategies. To ensure a fair comparison, we fix the sampling budget at 32 frames and disable the adaptive keyframe retrieval mechanism in Venus. Table II reports the reasoning accuracy and average query response latency under different settings. Venus consistently achieves comparable or even better accuracy than the baselines, while significantly outperforming them in response latency. Specifically, in terms of accuracy, Venus achieves performance comparable to AKS and BOLT, while outperforming the vanilla edge–cloud disaggregated architecture. This improvement is attributed to our sampling-based frame retrieval strategy, which is designed to ensure both relevance and diversity in keyframe selection. In terms of latency, Venus reduces redundant frames through scene segmentation and frame clustering, requiring embedding only a highly sparse set of indexed frames. This enables real-time perception and memory construction on edge devices. Under cloud-only deployment, BOLT and AKS require uploading the entire query video clip, resulting in significant transmission overhead. Venus achieves up to $9.9\times$ speedup on the Video-MME Short dataset. The performance gain increases with the length of the queried clip and reaches up to $126\times$ on the Video-MME Long subset. While edge–cloud deployment can reduce the volume of transmitted data, the frame-wise processing required by BOLT and AKS introduces considerable on-device computation overhead. This results in a response latency of up to $418.9$ seconds on the Video-MME Short dataset, and the latency continues to increase as the length of the queried video clip grows.
@@ V-E Ablation Study on Adaptive Keyframe Retrieval
In this section, we evaluate the impact of our Adaptive Keyframe Retrieval (AKR) module. We compare Venus with AKR ($N_{\text{max}}$=32) against Venus using sampling-based frame retrieval with fixed sampling budgets of 64 and 32 frames, evaluating reasoning accuracy, as well as average inference latency and communication overhead. We conduct the evaluation on the Video-MME Short dataset. To further highlight the advantage of AKR, we used ChatGPT-4o to help identify 60 queries (Video-MME subset) that focus on specific scenes within the video and therefore require only a small number of frames for accurate reasoning, similar to the example shown in Figure 9(left). The evaluation results are shown in Figure 11. We observe that Venus with AKR achieves accuracy comparable to fixed-budget baselines with 64 or 32 frames on Video-MME, while selecting only about 17 frames on average. This leads to a $1.6\times$$3.3\times$ reduction in VLM inference and communication overhead. On the curated subset, AKR further amplifies its advantage, achieving a $3.8\times$$7.6\times$ reduction in latency. By eliminating redundant frames that may interfere with VLM inference, AKR also outperforms the fixed-budget baselines in reasoning accuracy.
@@ VI RELATED WORK
Pre-VLM Era Online Video Understanding. Prior to the VLM era, CNNs [he2016deep, terven2023comprehensive] served as the backbone of online video understanding. Several studies introduce system-level optimizations for online video analysis. Focus [hsieh2018focus] supports low-cost ingest-time analytics on live video that later facilitates low-latency queries. Liu et al. [liu2019edge] design an online system that enables high-accuracy object detection at 60 FPS. Reducto [li2020reducto] designs an on-camera filtering module to enable real-time video analytics pipelines.
Video Understanding with VLMs VLMs’ visual-linguistic reasoning has driven their broad use in video understanding. LLaVA-OneVision [li2024llava] presents a family of open VLMs for multi-image and video-based visual tasks. AKS [tang2025adaptive] employs an adaptive optimization algorithm to enhance the coverage of selected keyframes. BOLT [liu2025bolt] introduces an adaptive frame selection strategy based on inverse transform sampling. Goldfish [ataallah2024goldfish] design a methodology tailored for comprehending videos of arbitrary lengths.
Retrieval-Augmented Generation for Video Understanding Recent works apply RAG [gao2023retrieval] to VLMs for long video understanding. VideoRAG [luo2024video] retrieves three types of visually-aligned auxiliary texts to boost VLM performance. DrVideo [ma2024drvideo] reformulates video understanding as long-document reasoning by converting videos into textual form. Ren et al. [ren2025videorag] introduce graph-based textual grounding to capture cross-video semantic relationships for long video understanding.

### v2

@@ V-A3 Baselines Methods
    MDF [9] proposes a query-irrelevant method that adopts a self-adaptive strategy to filter the most dominant frames.
@@
    Video-RAG [24] adopts uniform sampling over video frames and incorporates additional auxiliary data to construct a RAG database that enhances VLM performance.
@@
    AKS [34] proposes an adaptive keyframe selection method and employs an optimization algorithm to ensure comprehensive coverage of the selected keyframes.
@@
    BOLT [23] designs an inverse transform sampling method to prioritize query-relevant frames, improving performance on multi-source retrieval VQA tasks.
@@ V-C Comparison with Query-Relevant Baselines
We evaluate Venus against query-relevant baselines across various datasets, VLMs, and deployment strategies. For fairness, we fix the sampling budget to 32 frames and disable Venus’s adaptive keyframe retrieval. Table II shows that Venus matches or exceeds baseline accuracy while significantly reducing latency. Specifically, our sampling-based retrieval ensures relevance and diversity, achieving parity with AKS and BOLT while outperforming vanilla edge–cloud disaggregated architectures. Regarding efficiency, Venus’s scene segmentation and clustering minimize indexed frames, enabling real-time edge perception. In cloud-only setups, Venus achieves up to $126\times$ speedup on Video-MME Long by avoiding the massive transmission overheads inherent in BOLT and AKS. In edge–cloud deployments, Venus avoids the high on-device computation costs of frame-wise processing, which can result in latencies up to 418.9 seconds on Video-MME Short and scales further with video length.
@@ V-E Ablation Study on Adaptive Keyframe Retrieval
We compare Venus with AKR ($N_{\text{max}}$=32) against Venus using sampling-based frame retrieval with fixed sampling budgets of 64 and 32 frames, evaluating reasoning accuracy, as well as average inference latency and communication overhead. We conduct the evaluation on the Video-MME Short dataset. To further highlight the advantage of AKR, we used ChatGPT-4o to help identify 60 queries (Video-MME subset) that focus on specific scenes within the video and therefore require only a small number of frames for accurate reasoning, similar to the example shown in Figure 9(left). The evaluation results are shown in Figure 11. We observe that Venus with AKR achieves accuracy comparable to fixed-budget baselines with 64 or 32 frames on Video-MME, while selecting only about 17 frames on average. This leads to a $1.6\times$$3.3\times$ reduction in VLM inference and communication overhead. On the curated subset, AKR further amplifies its advantage, achieving a $3.8\times$$7.6\times$ reduction in latency. By eliminating redundant frames that may interfere with VLM inference, AKR also outperforms the fixed-budget baselines in accuracy.
@@ VI RELATED WORK
Pre-VLM Era Online Video Understanding. Focus [10] supports low-cost ingest-time analytics on live video that later facilitates low-latency queries. Liu et al. [22] design an online system that enables high-accuracy object detection at 60 FPS. Reducto [19] designs an on-camera filtering module to enable real-time video analytics pipelines.
Video Understanding with VLMs. LLaVA-OneVision [17] presents a family of open VLMs for video-based visual tasks. AKS [34] employs an adaptive optimization algorithm to enhance the coverage of selected keyframes. BOLT [23] introduces an adaptive frame selection strategy based on inverse transform sampling. Goldfish [1] design a methodology tailored for comprehending videos of arbitrary lengths.
Retrieval-Augmented Generation for Video Understanding. Recent works apply RAG [7] to VLMs for long video understanding. VideoRAG [24] retrieves three types of visually-aligned auxiliary texts to boost VLM performance. DrVideo [25] reformulates video understanding as long-document reasoning by converting videos into textual form. Ren et al. [32] introduce graph-based textual grounding to capture cross-video semantic relationships for long video understanding.
@@ VII Conclusion
Acknowledgments
This work was supported in part by the National Science Foundation of China (No. U25B2002); Guangdong S&T Programme (Grant No. 2024B0101040007); Guangdong Basic and Applied Basic Research Foundation (No. 2023B1515120058); Guangzhou Basic and Applied Basic Research Program (No. 2024A04J6367); The Program for Guangdong Introducing Innovative and Entrepreneurial Teams (No. 2017ZT07X355). National Natural Science Foundation of China (Grant No. 623B2093).